OpenStudy (anonymous):

the sum of two numbers is 36. their difference is 24 HELP AND EXPLAIN PLS

OpenStudy (radar):

x+y=36 x-y=24 -----------Add 2x=60 then x = 30 go ahead and solve for y

OpenStudy (anonymous):

y=6?

OpenStudy (radar):

Take one of the original like x-y=24 30-y=24 then y must be 6

OpenStudy (anonymous):

i was using subsitution wasn't working for me

OpenStudy (radar):

Yes very good and correct. I used elimination by adding eliminating the y. Substitution should of worked also.

OpenStudy (radar):

do you want to see how substitution would of worked?

OpenStudy (anonymous):

I understand I thought subsitution was the only way and then you brought up addition and I remebered thank you very much.

OpenStudy (radar):

No problem.

OpenStudy (anonymous):

one more quesrion: The sum of two numbers is 10. Three times the larger decreased by twice the smaller is 15. What are the numbers